# Retrieval-Augmented Generation in Biomedicine: A Survey of Technologies, Datasets, and Clinical Applications

### v2

@@ Introduction
Recent literature highlights several innovative RAG applications in clinical research leveraging different retrieval approaches. PodGPT 34 implements the BAAI embedding model for scientific literature retrieval informing research design and methodology development. LITURAt 62 utilizes the Entrez API for PubMed retrieval, supporting scientific data analysis in clinical research through relevant literature identification. These applications showcase RAG’s ability to optimize and improve the clinical research process through intelligent knowledge retrieval, highlighting its potential as a valuable tool for both research and education, and ultimately accelerating the translation of research findings into clinical practice.
@@
Recent research indicates that RAG systems can be both maliciously and unintentionally manipulated to produce unreliable decisions and potentially harm humans 1, 231. Such vulnerabilities are particularly unacceptable in the biomedical domain due to the potential for severe consequences.
Several dimensions of trustworthiness in biomedical RAG systems raise critical concerns. Robustness refers to the system’s ability to exclude malicious content introduced by attackers or filter inadvertent misinformation contained in web corpora that might contaminate the knowledge source 181. Although certain re-ranking methodologies described in Section Reranker can partially mitigate these issues, significant challenges remain in biomedical RAG applications. Explainability requires a comprehensive understanding of user intent to generate reliable responses aligned with ground-truth documents. A notable challenge emerges from diverse linguistic styles in input information, creating misalignment between user queries and knowledge sources. For instance, in medical QA systems, users may formulate queries in conversational language while knowledge sources are typically based on academic publications 139. This linguistic discrepancy poses substantial challenges for biomedical information retrieval, consequently diminishing the performance of biomedical RAG systems. Fairness dictates that trustworthy RAG systems must avoid discriminatory outputs during the generation process. Due to the uneven distribution of demographic attributes in training corpora 2, which may be present in both the pre-training data of LLMs and the knowledge sources utilized in RAG systems, biomedical RAG applications are susceptible to intrinsic biases. Addressing the fairness concerns also represents an important direction for future research.
@@
Among our surveyed publications, English remains the dominant language in all Biomedical RAG systems. This predominance is primarily attributed to the dominance of English in public biomedical knowledge repositories and medical task datasets. The performance of biomedical RAG systems in other languages, particularly those with limited resources, remains inadequately explored 232, 233. This fact presents two significant challenges in the development of multi-lingual biomedical RAG systems.
First, training LLMs specialized in low-resource languages is inherently difficult, which diminishes the generative capabilities of Biomedical RAG in other linguistic contexts. Second, the scarcity of available knowledge sources in non-dominant languages significantly constrains the augmentation potential afforded by knowledge retrieval mechanisms. Although some researchers have attempted to address this limitation through synthetic data generated by LLMs or translated from resource-rich languages 234, consistent and sustained efforts are still required to overcome this challenge in multilingual biomedical information retrieval and generation.
@@
Another challenge lies in extending retrieval capabilities beyond textual data to encompass multi-modalities highly prevalent in the biomedical domain, including images (e.g. X-ray 203, 235, 236), time series (e.g. ECG 152), sound recordings (e.g. lung sounds recorded 237), and video 238. Among the studies we identified, the overwhelming majority focus exclusively on text-based retrieval, with only a limited subset addressing image-based retrieval, while other modalities remain largely unexplored. Multi-modal data presents several intractable challenges for retrieval systems. For instance, in medical imaging, although we can utilize a visual encoder model as a retriever, high intra-class variance significantly complicates image retrieval processes 239. Specifically, identical medical conditions can manifest with remarkable heterogeneity across visual representations; for instance, images of histologically identical tumors may exhibit substantially different morphological characteristics when visualized across different patients or at varying stages of disease progression. Attention also should be paid to developing cross-modal RAG in the Biomedical domain and improving its accuracy and trustworthiness in the future.
@@
The application of RAG in the biomedical domain faces significant limitations due to computing resource constraints, which restrict its utility in practical scenarios. Within the RAG framework, LLMs play a crucial role in contextual understanding and human-like response generation based on retrieved knowledge 1. However, these models typically require high-performance computing clusters to operate effectively. For instance, the minimum deployment unit of DeepSeek V3’s decoding stage consists of 40 nodes with 320 GPUs 240.
Empirical research consistently demonstrates that LLMs with larger parameter counts generally exhibit superior capacity for understanding and generation tasks 241, 242. Nevertheless, in real-world healthcare environments such as hospitals, maintaining high-performance computing infrastructure is often impractical and cost-prohibitive. Several promising approaches have emerged to address these constraints. One solution involves developing smaller, distilled LLMs derived from their larger counterparts 243. Models such as QWQ 160 and Mistral Small 3.1 40 demonstrate competitive performance despite having significantly reduced parameter counts. Applying Portable LLMs 244 as generation models is also a promising method. Another viable approach incorporates RAG systems that combine local knowledge bases with external LLM API services. This hybrid configuration presents an acceptable compromise when computing resources are limited, although it introduces potential privacy concerns, as local data must be transmitted to commercial providers for response generation.
@@
Privacy concerns in biomedical RAG systems exist throughout both retrieval and generation processes. During knowledge retrieval, reference documents frequently contain patients’ clinical histories, which may include personally identifiable information that individuals would prefer to keep confidential 245. Such information handling might not comply with stringent regulatory frameworks such as the General Data Protection Regulation 246 and the Health Insurance Portability and Accountability Act 247. Furthermore, the generation phase presents additional challenges, particularly when utilizing commercial LLMs for biomedical RAG development. This approach necessitates the transmission of sensitive data across the internet, significantly increasing the risk of data breaches and unauthorized access. Consequently, developing biomedical RAG systems with robust privacy-preserving mechanisms remains a critical direction for future research endeavors.

### v3

@@ Introduction
Recent literature highlights several innovative RAG applications in clinical research leveraging different retrieval approaches. PodGPT jia2024podgpt implements the BAAI embedding model for scientific literature retrieval informing research design and methodology development. LITURAt Peasley2025 utilizes the Entrez API for PubMed retrieval, supporting scientific data analysis in clinical research through relevant literature identification. These applications showcase RAG’s ability to optimize and improve the clinical research process through intelligent knowledge retrieval, highlighting its potential as a valuable tool for both research and education, and supporting the translation of research findings into clinical practice.
Discussion
In this section, we summarize the key observations drawn from technological approaches, datasets, evaluation resources, and biomedical applications, examines the limitations of existing work, and further discusses the major challenges faced by RAG systems in the biomedicine domain while outlining promising directions for future research and development.
Summary of technological approaches
In Section.III, we reviewed three key technological foundations of biomedical RAG systems identified in our survey: retrieval, reranking, and generation. Retrieval methods span traditional sparse models such as BM25bm25, multiple classes of dense retrievers—including general PLMs, commercial embedding APIs, domain-adapted biomedical models, LLM-based embeddings, and specialized architectures such as MedCPTjin2023medcpt and BMRetrieverxu2024bmretriever, and multimodal frameworks. Knowledge-graph–augmented retrieval further enhances precision by leveraging structured biomedical ontologies and explicit entity–relation pathways. Benefiting from its efficiency and proven generalizability, BM25bm25 remains the leading approach for retrieval; however, it is limited to textual resources. Emerging image-based and multimodal systems remain limited, partly because single-vector indexingXia2024RULE, pmlr-v219-ranjit23a, Sun2024FactAware—widely used in text retrieval—may underrepresent the heterogeneous characteristics of multimodal data, making their indexing and retrieval methodologies for multimodal RAG still worthy of further exploration.
Reranking techniques provide a second-stage selection mechanism, aiming to enhance the quality of the initial rank with more advanced and algorithmically complex strategies. Existing systems employ statistical approaches (e.g., RRF, MMR), transformer-basedjia2024podgpt, adejumo2024retrieval and cross-encoder modelsjin2023medcpt, LLM-based relevance classifiersdarnell2024biomedical, craig2024lmrac, as well as domain-specific rerankers designed to better capture biomedical semantics. The main challenge for reranking models is generalization at inference time, as they often rely on annotated samples. Creating application-specific rank–gold labels for existing datasets is expensive, while automatic annotation carries risks of reduced accuracy, forcing developers to carefully balance this trade-off in practice.
Lastly, on the generation side, both open-source and commercial LLMs are used extensively, alongside an increasing number of biomedical-specialized models (see Table LABEL:tab:llm_taxonomy), to generate content benefiting from the retrieved biomedical resources. We also find that domain adaptation—whether applied to retrievers, rerankers, or generators—enhances terminology handling, conceptual grounding, and factual reliability. Building on these observations, commercial LLMs are easy to use and can deliver competitive performance, whereas open-source models—especially biomedical-specialized ones—allow continued training or fine-tuning for specific tasks and even modifications to the model architecture, enabling more flexible application across diverse tasks. However, this flexibility also imposes additional computational resource demands on developers.
For the future development of RAG, we believe that continued progress in multimodal retrieval and knowledge graph integration will expand the scope of information accessible to RAG systems and improve their ability to generate contextually informed, clinically aligned outputs. This underscores the growing importance of cross-modal fusion strategies and targeted domain adaptation in advancing biomedical RAG capabilities.
Summary of dataset and evaluation approaches
In Sec.IV, we reviewed datasets relevant to the development and evaluation of biomedical RAG systems, categorizing them into Knowledge Source Datasets (Table LABEL:tab:med_knowledge) and Medical Task Datasets (Table LABEL:tab:biomedical-comprehensive). Biomedical knowledge sources are highly diverse, and this heterogeneity poses significant challenges for retrieval, as different sources often exhibit distinct linguistic styles and vocabulary patterns.
On the evaluation side, current benchmarks mainly rely on various Medical Task Datasets, with a particular focus on knowledge-intensive tasks such as Multiple-Choice Examinations (e.g., MedQA jin2020disease, MedMCQA pal2022medmcqa), which simulate medical licensing assessments, and a range of Question Answering (QA) tasksBenAbacha-BMC-2019, singhal2022large, hou2023geneturing, Lozano2024ClinfoAI, ALONSO2024102938. However, existing evaluations still lack tasks that are more concrete and aligned with real-world clinical scenarios. We also observe a strong dominance of English-language datasets and a heavy reliance on textual benchmarks. Substantial future efforts are needed to build richer cross-lingual and multi-modal evaluation resources that better reflect the diversity and complexity of biomedical practice.
For evaluation metrics, we follow the retrieval–reranking–generation pipeline and review the metrics used at each stage accordingly. Retrieval and reranking are typically assessed using standard information-retrieval measures such as recall, precision, nDCG, and MRR, while generation metrics vary widely depending on the specific downstream task. To provide a clearer overview, we group generation-level evaluations into four major task types, each associated with its own metric family. Although many automatic metrics exist, they often struggle to capture the nuanced reasoning quality required in complex biomedical dialogue and decision-support scenarios metric-survey. As a result, human evaluation remains essential. However, due to the substantial cost of expert annotation, LLM-based evaluators are increasingly adopted Croxford2025, metric-survey. These models can follow instructions and emulate human evaluation criteria, thereby offering a more scalable alternative to manual assessment, although they may also introduce risks such as hallucination and evaluation bias. Therefore, developing evaluation frameworks that are both low-cost and highly accurate remains an important direction for future research.
Summary of RAG applications in biomedicine
Biomedical RAG applications (Sec. IV) span a wide range of healthcare domains—from clinical decision support and precision medicine to medical education and clinical research—highlighting RAG’s capacity to enhance factual grounding, interpretability, and domain specificity for AI-driven healthcare solutions. Across the surveyed literature, RAG-based systems have demonstrated utility in various medical specialties, including oncologyLammert2024ExpertLLM, radiologysloan2024automated, ophthalmologychen2024eyegpt, geneticsli2024favorgpt, Lu2025BoostingGPT, Fukushima2025, nursingZHAO2025104284, and rare disease diagnosisZHAO2025104284, Zelin2024RareDisease, garcia2024deepphenotyping. These domain-specific implementations suggest that retrieval-augmented approaches can help connect LLMs with biomedical evidence, thereby supporting grounded question answering, enhancing the consistency of clinical reasoning processes, and contributing to clearer and more interpretable outputs for both patient-facing and clinician-facing use cases. However, the surveyed applications also reveal several practical constraints. Retrieval steps often introduce latency and increase token length, which can reduce system efficiency and limit scalability in real clinical workflowsAmmar2025. In addition, current retrievers still struggle with fine-grained biomedical semantics—such as distinguishing closely related clinical entities or integrating cross-modal evidence—which can lead to suboptimal grounding even when high-quality sources existTan2024. Addressing these issues will be important for deploying RAG systems in settings that demand both timely responses and high diagnostic precision.
@@ Discussion
Recent research indicates that RAG systems can be both maliciously and unintentionally manipulated to produce unreliable decisions and potentially harm humans Fan2024RAG, deng2024pandora. Such vulnerabilities are particularly unacceptable in the biomedical domain due to the potential for severe consequences.
Several dimensions of trustworthiness in biomedical RAG systems raise critical concerns. Robustness refers to the system’s ability to exclude malicious content introduced by attackers or filter inadvertent misinformation contained in web corpora that might contaminate the knowledge source gao2023retrieval. Although certain re-ranking methodologies described in Section 3.2 can partially mitigate these issues, significant challenges remain in biomedical RAG applications. Explainability requires a comprehensive understanding of user intent to generate reliable responses aligned with ground-truth documents. A notable challenge emerges from diverse linguistic styles in input information, creating misalignment between user queries and knowledge sources. For instance, in medical QA systems, users may formulate queries in conversational language while knowledge sources are typically based on academic publications lee2024enhancing. This linguistic discrepancy poses substantial challenges for biomedical information retrieval, consequently diminishing the performance of biomedical RAG systems. Fairness dictates that trustworthy RAG systems must avoid discriminatory outputs during the generation process. Due to the uneven distribution of demographic attributes in training corpora he2023survey, which may be present in both the pre-training data of LLMs and the knowledge sources utilized in RAG systems, biomedical RAG applications are susceptible to intrinsic biases. Addressing the fairness concerns also represents an important direction for future research.
@@
Among our surveyed publications, English remains the dominant language in all Biomedical RAG systems. This predominance is primarily attributed to the dominance of English in public biomedical knowledge repositories and medical task datasets. The performance of biomedical RAG systems in other languages, particularly those with limited resources, remains inadequately explored neveol2018clinical, Englishllm. This fact presents two significant challenges in the development of multi-lingual biomedical RAG systems.
First, training LLMs specialized in low-resource languages is inherently difficult, which diminishes the generative capabilities of Biomedical RAG in other linguistic contexts. Second, the scarcity of available knowledge sources in non-dominant languages significantly constrains the augmentation potential afforded by knowledge retrieval mechanisms. Although some researchers have attempted to address this limitation through synthetic data generated by LLMs or translated from resource-rich languages mondshine2025beyond, consistent and sustained efforts are still required to overcome this challenge in multilingual biomedical information retrieval and generation.
@@
Another challenge lies in extending retrieval capabilities beyond textual data to encompass multi-modalities highly prevalent in the biomedical domain, including images (e.g. X-ray johnson2019mimic, pelka2018roco, lin2023pmc), time series (e.g. ECG pmlr-v225-yu23b), sound recordings (e.g. lung sounds recorded fraiwan2021dataset), and video gupta2023dataset. Among the studies we identified, the overwhelming majority focus exclusively on text-based retrieval, with only a limited subset addressing image-based retrieval, while other modalities remain largely unexplored. Multi-modal data presents several intractable challenges for retrieval systems. For instance, in medical imaging, although we can utilize a visual encoder model as a retriever, high intra-class variance significantly complicates image retrieval processes zhong2024class. Specifically, identical medical conditions can manifest with remarkable heterogeneity across visual representations; for instance, images of histologically identical tumors may exhibit substantially different morphological characteristics when visualized across different patients or at varying stages of disease progression. Attention also should be paid to developing cross-modal RAG in the biomedical domain and improving its accuracy and trustworthiness in the future.
@@
The application of RAG in the biomedical domain faces significant limitations due to computing resource constraints, which restrict its utility in practical scenarios. Within the RAG framework, LLMs play a crucial role in contextual understanding and human-like response generation based on retrieved knowledge Fan2024RAG. However, these models typically require high-performance computing clusters to operate effectively. For instance, the minimum deployment unit of DeepSeek V3’s decoding stage consists of 40 nodes with 320 GPUs deepseekv3technicalreport.
Empirical research consistently demonstrates that LLMs with larger parameter counts generally exhibit superior capacity for understanding and generation tasks brown2020language, hoffmann2022training. Nevertheless, in real-world healthcare environments such as hospitals, maintaining high-performance computing infrastructure is often impractical and cost-prohibitive. Several promising approaches have emerged to address these constraints. One solution involves developing smaller, distilled LLMs derived from their larger counterparts xu2024survey. Models such as QWQ qwen32b_2024 and Mistral Small 3.1 mistral_small_24b_2024 demonstrate competitive performance despite having significantly reduced parameter counts. Applying portable LLMs khan2024portllm as generation models is also a promising method. Another viable approach incorporates RAG systems that combine local knowledge bases with external LLM API services. This hybrid configuration presents an acceptable compromise when computing resources are limited, although it introduces potential privacy concerns, as local data must be transmitted to commercial providers for response generation.
@@
Privacy concerns in biomedical RAG systems exist throughout both retrieval and generation processes. During knowledge retrieval, reference documents frequently contain patients’ clinical histories, which may include personally identifiable information that individuals would prefer to keep confidential meystre2010automatic. Such information handling might not comply with stringent regulatory frameworks such as the General Data Protection Regulation GDPR2016 and the Health Insurance Portability and Accountability Act HIPAA1996. Furthermore, the generation phase presents additional challenges, particularly when utilizing commercial LLMs for biomedical RAG development. This approach necessitates the transmission of sensitive data across the internet, significantly increasing the risk of data breaches and unauthorized access. Consequently, developing biomedical RAG systems with robust privacy-preserving mechanisms remains a critical direction for future research endeavors.